Factor the quadratic expression completely, and find the roots of the expression.
6x^2 - 42x + 72

Answer by checkley71(8403)

6X~2-42X+72=0 OR DIVIDING ALL FACTORS BY 6 WE GET X~2-7X+12 OR (X-4)(X-3)=0 OR
X-4=0 OR X=4 & X-3=0 OR X=3
PROOF 6*4~2-42*4+72=0 OR 6*16-168+72=0 OR 96-168+72=0 OR 168-168=0 OR 0=0
6*3~2-42*3+72=0 OR 6*9-126+72=0 OR 54-126+72=0 OR 126-126=0 OR 0=0